Circular No. 7761

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001gd IN NGC 5033 S. Nakano, Sumoto, Japan, reports the discovery of an apparent su-

pernova (mag 16.5) by K. Itagaki, Yamagata, Japan, on CCD images taken with a 0.60-m f/6.0 reflector on Nov. 24.820 UT. Itagaki remarked that the new object is located ∼ 3′ north-northwest of the center of NGC 5033, and he added that nothing was visible at this position on a CCD frame taken on May 29 (limiting mag 19). Nakano forwards the following precise position for SN 2001gd, which is the average of measurements by Y. Kushida from two CCD images taken by R. Kushida, Yatugatake South Base Observatory, with a 0.40-m f/6.3 Schmidt-Cassegrain reflector (+ CCD) in bright twilight on Nov. 24.851 (when the new object was at mag 17.5) and 25.802 (mag 16.2): α = 13^{h}13^{m}23^{s}.89, δ = +36^{o}38′17′′.7 (equinox 2000.0), which is 3′23′′ in p.a. 347^{o} from the nucleus of NGC 5033. R. Kushida adds that there is no star visible at this location on her earlier frames taken from 1996 to Apr. 2001. SNe 1950C and 1985L appeared in this galaxy. A. Dimai, Cortina d’Ampezzo, Italy, reports his independent discovery of SN 2001gd (mag ∼ 16.4) on images taken in the course of the CROSS program (cf. IAUC 7373) with a 0.51-m telescope on Nov. 25.21 and 26.21; his position end figures for the new object are 23^{s}.92, 18′′.2, which is ∼ 52′′ west and 161′′ north of the nucleus of NGC 5033. A CROSS image taken on May 8 showed nothing at this position (limiting mag ∼ 18.5), and SN 2001gd does not appear on Palomar Sky Survey red and blue plates.

SUPERNOVAE 2001gb AND 2001gc T. Matheson, S. Jha, P. Challis, and R. Kirshner, Harvard-Smithsonian

Center for Astrophysics, report that spectra (range 370–750 nm) were obtained of SN 2001gb (cf. IAUC 7758) and SN 2001gc (cf. IAUC 7759) on Nov. 22.49 and 22.25 UT by M. Calkins with the Mt. Hopkins 1.5-m telescope (+ FAST spectrograph), showing them both to be type-Ia supernovae. Adopting the NASA/IPAC Extragalactic Database (NED) recession velocity of 7714 km/s for the host galaxy of SN 2001gb, the supernova expansion velocity is ∼ 11 300 km/s for Si II (rest 635.5 nm); the spectral-feature age of SN 2001gb (Riess et al. 1997, A.J. 114, 722) is 0 ± 2 days after maximum light. Adopting the NED recession velocity of 5783 km/s for the host galaxy, the expansion velocity of SN 2001gc is ∼ 16 600 km/s for Si II (rest 635.5 nm); this expansion velocity implies an age of ≈ 2 weeks before maximum, based upon comparison with values presented by Jha et al. (1999, Ap.J. Supp. 125, 73).

2001 November 26  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green